# Decentralized Credential Status Management: A Paradigm Shift in Digital Trust

### v1

@@ Abstract
Public key infrastructures are essential for Internet security, ensuring robust certificate management and revocation mechanisms. The transition from centralized to decentralized systems presents challenges such as trust distribution and privacy-preserving credential management. The transition from centralized to decentralized systems is motivated by addressing the single points of failure inherent in centralized systems and leveraging decentralized technologies’ transparency and resilience. This paper explores the evolution of certificate status management from centralized to decentralized frameworks, focusing on blockchain technology and advanced cryptography. We provide a taxonomy of the challenges of centralized systems and discuss opportunities provided by existing decentralized technologies. Our findings reveal that, although blockchain technologies enhance security and trust distribution, they represent a bottleneck for parallel computation and face inefficiencies in cryptographic computations. For this reason, we propose a framework of decentralized technology components that addresses such shortcomings to advance the paradigm shift toward decentralized credential status management.
@@ I Introduction
Our research contributes to the management of trust and credential status in decentralized networks. We provide insights into managing digital trust and credentials by introducing a taxonomy that outlines the limitations of CPKI and explore opportunities within DPKI systems. The research questions addressed are: (1) How can decentralized technologies enhance credential status management? (2) What cryptographic primitives and governance frameworks are required for robust DCSM? This paper offers a detailed comparative analysis of traditional PKI components and decentralized approaches. We summarize the findings as a technology framework that serves as a foundation for future research and innovation in this critical area.
@@ III Related Work
This section reviews literature on decentralized PKI and self-sovereign identity.
@@
Similarly, CertLedger is a blockchain-based system for transparent and secure certificate management and revocation proposed by Kubilay et al. [27]. CertLedger focuses on removing single points of failure associated with CAs and ensuring the integrity of the certificate lifecycle through a decentralized ledger. However, like Certcoin and PB-PKI, CertLedger does not incorporate SSI principles or advanced cryptographic techniques such as MPC and FHE, limiting its ability to provide user-centric and privacy-preserving identity management solutions.
Conversely, Mühle et al. [28] provide a comprehensive SSI survey that outlines its principles, benefits, and challenges. SSI emphasizes user control over digital identities through DIDs and VC, enabling users to manage their credentials without relying on centralized authorities. The survey emphasizes privacy and user autonomy, but lacks technical details on integrating SSI with DPKI. Furthermore, the survey does not address the use of advanced cryptographic techniques such as MPC and FHE.
@@ V-F Decentralized Credential Status Management (DCSM)
is a cryptographic technique that enables multiple parties to collaboratively compute a function over their private inputs without revealing the underlying data. In the context of DCSM, integrating MPC facilitates secure and distributed credential issuance and (un)revocation processes, ensuring that no single entity has complete control over the credential lifecycle. MPC is particularly relevant in DCSM, where trust is decentralized and no single entity has complete visibility and control over the processed data. However, the practical deployment of MPC in DCSM faces computational and communication challenges due to its complexity. Ongoing research focuses on improving the efficiency and scalability of MPC protocols to enable seamless integration with high-throughput blockchain environments, bridging the gap between theoretical guarantees and practical applications (R1, R2, R6, R7) [67].
@@
is an encryption scheme that allows computations on encrypted data without decryption. FHE is crucial in preserving the privacy of sensitive credential data throughout its lifecycle, from issuance and storage to verification and (un)revocation. By leveraging FHE, DCSM implementations can ensure that credential status data remains encrypted while allowing necessary computations and updates without requiring any party to access the credentials’ status raw data. Combining FHE with MPC, DCSM implementations can enable secure and private computation of encrypted credential data across multiple parties, ensuring that no single entity can access the raw data. Thus, the combination of MPC and FHE provides a privacy-preserving toolset for decentralized systems, enabling distributed credential issuance, verification, and (un)revocation. However, the practical implementation of FHE in DCSM implementations faces challenges primarily related to computational complexity. Ongoing research efforts aim to optimize FHE algorithms, develop more efficient implementations, and explore partially homomorphic encryption (PHE) schemes to mitigate these challenges (R1, R2, R4, R6) [68].
@@
are cryptographic primitives that allow the verification of claims without revealing additional information beyond the validity of the claim. In DCSM, ZKP plays a cruical role in enabling privacy-preserving credential verification, allowing users to prove possession of a credential without disclosing sensitive information. Integrating ZKP with blockchain technology ensures the security and immutability of the verification process, as proofs are recorded on a tamper-proof ledger. However, the practical implementation of ZKP in DCSM faces proof and computational efficiency challenges. Ongoing research efforts aim to optimize proof construction and verification times, reduce the on-chain computation load of proofs, and improve the user-friendliness of ZKP-based systems. Exploring novel cryptographic primitives, such as zk-SNARKs, zk-STARKs, and ZoKrates, which remove the need for a trusted setup, holds promise for the future of ZKPs in DCSM (R1, R3, R4, R5) [69].
@@
provides robust security with shorter key lengths than traditional cryptographic systems, such as RSA. Shorter key lengths enhance the processing speed and reduce computational overhead, making ECC suitable for DCSM applications, especially on devices with limited resources. Using ECC, DCSM can improve the cryptographic security of credential transactions while ensuring efficient operational performance across the decentralized network (R1, R2, R4, R7) [70].
@@ VII Future Work
Future research should prioritize optimizing the computational efficiency of ZKP, MPC, and FHE to improve scalability and performance. Emerging technologies such as sharding [74] and Layer 2 [75] offer promising avenues to overcome scalability limitations in blockchain-based DCSM implementations. Developing standardized protocols and interfaces is crucial for integrating these technologies with SSI frameworks. In addition, practical deployment studies are needed to explore real-world applications and address regulatory challenges associated with decentralized systems. Focusing on these areas can improve the robustness and practicality of DCSM systems, meeting the evolving needs of digital credential management.

### v2

@@ Abstract
Public key infrastructures are essential for Internet security, ensuring robust certificate management and revocation mechanisms. The transition from centralized to decentralized systems presents challenges such as trust distribution and privacy-preserving credential management. The shift to decentralized systems addresses single points of failure in centralized systems and leverages the transparency and resilience of decentralized technologies. This paper explores the evolution of certificate status management from centralized to decentralized frameworks, focusing on blockchain technology and advanced cryptography. We provide a taxonomy of the challenges of centralized systems and discuss opportunities provided by existing decentralized technologies. Our findings reveal that, although blockchain technologies enhance security and trust distribution, they represent a bottleneck for parallel computation and face inefficiencies in cryptographic computations. For this reason, we propose a framework of decentralized technology components that addresses such shortcomings to advance the paradigm shift toward decentralized credential status management.
@@ I Introduction
Trust, digital trust, and decentralized trust are foundational concepts in the security and reliability of digital systems, particularly within decentralized environments [4]. Trust refers to the confidence in the integrity, reliability, and security of systems and entities. Digital trust encompasses this confidence within digital interactions, relying on cryptographic technologies and protocols. Decentralized trust distributes the trust anchor among multiple nodes or entities, enhancing resilience and mitigating single points of failure.
@@
Our research contributes to the management of trust and credential status in decentralized networks. We provide insights into managing digital trust and credentials by introducing a taxonomy that outlines the limitations of CPKI and explore opportunities within DPKI systems. The research questions addressed are: (1) How can decentralized technologies enhance credential status management? (2) What cryptographic primitives and governance frameworks are required for robust DCSM? (3) How can DCSM be effectively integrated into the Self-Sovereign Identity (SSI) [6] paradigm to improve user-centric and privacy-preserving identity management?
This paper offers a detailed comparative analysis of traditional PKI components and decentralized approaches. We summarize the findings as a technology framework that serves as a foundation for future research and innovation in this critical area.
@@ III Related Work
This section provides a comprehensive overview of existing literature on decentralized credential management systems, focusing on their evolution, challenges, and technological advancements. The related work section examines DPKI models, highlighting their strengths, limitations, and privacy considerations. Additionally, this section reviews existing surveys on trust-based systems. This structured approach establishes a solid foundation for this paper’s subsequent discussion and analysis. Furthermire, this section reviews literature on decentralized PKI and Self-Sovereign Identity.
@@
Our proposed framework leverages decentralized technologies and advanced cryptographic techniques to enhance security, privacy, and efficiency in credential management. Unlike Certcoin, which links identities and public keys on a public ledger, our framework uses ZKP to allow users to prove possession of a credential without revealing their identity. This approach ensures that user identity remains private while providing verifiable proof of credential ownership. Additionally, MPC allows multiple parties to compute functions over their inputs without revealing the inputs themselves, enhancing privacy during credential issuance and verification. FHE enables computations on encrypted data, ensuring that sensitive information remains confidential. These techniques collectively address privacy concerns associated with publicly linking identities and public keys, providing a secure and privacy-preserving solution.
@@
Compared to PB-PKI, which addresses privacy by unlinking identities from public keys, our proposed framework further enhances privacy by integrating MPC and FHE. MPC enables multiple entities to collaboratively process credential information without exposing their private data, ensuring that sensitive details remain confidential. FHE allows computations on encrypted data, providing an additional layer of security and privacy. By combining MPC and FHE, our framework ensures that credential transactions are secure and private, mitigating risks associated with data exposure. A combined integration of MPC and FHE represents an advancement over PB-PKI, offering robust privacy protection while maintaining the integrity and verifiability of credentials.
The study by Vanin et al. [29] introduces an FHE approach to ensure end-to-end data protection in healthcare. The proposed system of Vanin et al. aims to mitigate privacy risks and enhance data interoperability among health institutions by utilizing distributed hash tables (DHT) and the InterPlanetary File System (IPFS) for data storage and sharing. The model’s effectiveness is evaluated by its ability to protect personal privacy while enabling encrypted data calculations without exposing sensitive information. Our work is distinct from the study by Vanin et al. as we concentrate explicitly on decentralized credential status management.
Similarly, CertLedger is a blockchain-based system for transparent and secure certificate management and revocation proposed by Kubilay et al. [30]. CertLedger focuses on removing single points of failure associated with CAs and ensuring the integrity of the certificate lifecycle through a decentralized ledger. However, like Certcoin, PB-PKI, and the research of Vanin et al., CertLedger does not incorporate SSI principles or advanced cryptographic techniques such as MPC and FHE, limiting its ability to provide user-centric and privacy-preserving identity management solutions.
Fragkos et al. [31] explore artificial intelligence approaches to enhance the performance and privacy of electronic cash systems. Fragkos et al. introduce PUF-Cash, an e-Cash scheme leveraging Physical Unclonable Functions (PUFs) for authentication and encryption between users, banks, and multiple trusted third parties. The AI methods utilized in PUF-Cash optimize the distribution of withdrawal amounts among trusted third parties to maximize privacy and performance. The scheme provided by Fragkos et al. aims to provide strong privacy guarantees, ensuring that users can conduct transactions anonymously, thereby preserving the anonymity implicit in traditional fiat money.
Comparison of CertLedger and PUF-Cash
CertLedger focuses on linking identity with public keys in a public ledger, ensuring transparency and security. However, it does not address privacy concerns, as identities and public keys are publicly linked, making it unsuitable for use cases requiring privacy. Unlike CertLedger, which emphasizes blockchain transparency over user-centric privacy mechanisms, Fragkos et al. leverages AI and PUFs to enhance privacy and performance in e-Cash transactions. PUF-Cash uses multiple trusted third parties and advanced cryptographic techniques to ensure the anonymity and security of electronic transactions.
Our proposed framework integrates SSI concepts to enhance user-centric and privacy-preserving identity management compared to CertLedger and PUF-Cash. SSI principles, including DIDs and VCs, grant users control over their digital identities without reliance on centralized authorities. Unlike CertLedger, which relies on blockchain transparency, and PUF-Cash, which employs AI for privacy, our approach leverages SSI principles to prevent unauthorized access and profiling, offering a more robust and privacy-focused solution for identity management.
Conversely, Mühle et al. [32] provide a comprehensive survey outlining SSI’s principles, benefits, and challenges. SSI emphasizes user control over digital identities through DIDs and VC, enabling users to manage their credentials without relying on centralized authorities. The survey emphasizes privacy and user autonomy but lacks technical details on integrating SSI with DPKI. Furthermore, the survey does not address using advanced cryptographic techniques such as MPC and FHE.
@@ IV-B Organizational Considerations
The GDPR presents specific challenges to blockchain-based systems due to its requirements for data protection and the right to be forgotten. Decentralized frameworks must ensure that personal data is handled in compliance with GDPR, which may involve designing mechanisms for data minimization and user consent. The European Union’s eIDAS [47] regulation provides a legal framework for electronic identities and trust services. Integrating eIDAS-compliant identity services within a decentralized framework promises to enhance legal recognition and trust across borders. Our proposed framework addresses legal complexities and enhances compliance by incorporating these regulatory considerations.
@@ V-F Decentralized Credential Status Management (DCSM)
is a cryptographic technique that enables multiple parties to collaboratively compute a function over their private inputs without revealing the underlying data. In the context of DCSM, integrating MPC facilitates secure and distributed credential issuance and (un)revocation processes, ensuring that no single entity has complete control over the credential lifecycle. MPC is particularly relevant in DCSM, where trust is decentralized and no single entity has complete visibility and control over the processed data. Furthermoe, MPC enhances trust in DCSM by distributing the processing of credential data among multiple parties, thus preventing any single point of control or failure. However, the practical deployment of MPC in DCSM faces computational and communication challenges due to its complexity. Ongoing research focuses on improving the efficiency and scalability of MPC protocols to enable seamless integration with high-throughput blockchain environments, bridging the gap between theoretical guarantees and practical applications (R1, R2, R6, R7) [72].
@@
is an encryption scheme that allows computations on encrypted data without decryption. FHE is crucial in preserving the privacy of sensitive credential data throughout its lifecycle, from issuance and storage to verification and (un)revocation. Privacy preservation with FHE ensures that credential data remains secure even during processing, thereby enhancing the overall security framework of DCSM. By leveraging FHE, DCSM implementations can ensure that credential status data remains encrypted while allowing necessary computations and updates without requiring any party to access the credentials’ status raw data. Combining FHE with MPC, DCSM implementations can enable secure and private computation of encrypted credential data across multiple parties, ensuring that no single entity can access the raw data. Thus, the combination of MPC and FHE provides a privacy-preserving toolset for decentralized systems, enabling distributed credential issuance, verification, and (un)revocation. However, the practical implementation of FHE in DCSM implementations faces challenges primarily related to computational complexity. Ongoing research efforts aim to optimize FHE algorithms, develop more efficient implementations, and explore partially homomorphic encryption (PHE) schemes to mitigate these challenges (R1, R2, R4, R6) [73].
@@
are cryptographic primitives that allow the verification of claims without revealing additional information beyond the validity of the claim. In DCSM, ZKP plays a cruical role in enabling privacy-preserving credential verification, allowing users to prove possession of a credential without disclosing sensitive information. ZKP enhances user trust by ensuring that only the necessary proof is shared, keeping the underlying data private. Integrating ZKP with blockchain technology ensures the security and immutability of the verification process, as proofs are recorded on a tamper-proof ledger. However, the practical implementation of ZKP in DCSM faces proof and computational efficiency challenges. Ongoing research efforts aim to optimize proof construction and verification times, reduce the on-chain computation load of proofs, and improve the user-friendliness of ZKP-based systems. Exploring novel cryptographic primitives, such as zk-SNARKs, zk-STARKs, and ZoKrates, which remove the need for a trusted setup, holds promise for the future of ZKPs in DCSM (R1, R3, R4, R5) [74].
@@
provides robust security with shorter key lengths than traditional cryptographic systems, such as RSA. Shorter key lengths enhance the processing speed and reduce computational overhead, making ECC suitable for DCSM applications, especially on devices with limited resources. ECC enables efficient cryptographic operations, which is critical for maintaining the performance of DCSM across diverse devices and environments. Using ECC, DCSM can improve the cryptographic security of credential transactions while ensuring efficient operational performance across the decentralized network (R1, R2, R4, R7) [75].
V-G Practical Implementation Examples
Several real-world scenarios can illustrate the practical implementation of DCSM. A systematic review by Agbo et al. [76] presents challenges and opportunities for decentralized healthcare applications to enhance the security and privacy of patient records. Patient records could be verifiable credentials, and our proposed DCSM framework could ensure secure management. By leveraging blockchain technology and advanced cryptographic techniques, healthcare providers can securely manage patient data and ensure access is restricted to authorized entities, complying with stringent privacy regulations like the Health Insurance Portability and Accountability Act (HIPAA) [77].
In the education sector, DCSM streamlines the verification of academic qualifications [78][79]. Universities can issue and verify verifiable credentials securely, enabling education institutions and employers to authenticate academic records without intermediaries. This process enhances verification efficiency and ensures the integrity and authenticity of educational records, fostering trust among stakeholders .
These examples demonstrate the proposed framework’s practical applicability and effectiveness in enhancing security, privacy, and trust across various domains.
@@ VI Discussion
Our framework employs several strategies and techniques to address scalability and performance challenges. Sharding divides on-chain and off-chain storage into smaller, more manageable segments, each capable of processing transactions independently. Sharding could improve the overall throughput and reduce latency [80]. Layer-2 solutions, such as state channels and sidechains, promise to improve the throughput of off-chain transactions, reducing the load on the main blockchain and increasing transaction speeds [81].
@@ VII Future Work
Future work will focus on optimizing the computational efficiency of ZKP, MPC, and FHE to improve scalability. Emerging technologies such as sharding [85] and Layer-2 [86] solutions offer promising avenues to overcome scalability limitations in blockchain-based DCSM implementations. Developing standardized protocols and interfaces is crucial for integrating these technologies with SSI frameworks. In addition, practical deployment studies are needed to explore real-world applications and address regulatory challenges associated with decentralized systems. Focusing on these areas can improve the robustness and practicality of DCSM systems, meeting the evolving needs of digital credential management.
Acknowledgment
This work was supported by the European Union’s Digital Europe Program (DIGITAL) research and innovation program under grant agreement number 101102743 (TRACE4EU). The authors thank the Trace4EU project partners for their invaluable contributions and the cheqd team for their support and expertise.